OpenStudy (anonymous):

volume of cylindrical shells (calculus): y^2=x, x=2y rotate about y axis

OpenStudy (anonymous):

bro whats the base equation again, something like pi int .... of the derivative squared?

OpenStudy (anonymous):

idk if that matters

OpenStudy (anonymous):

i have my integral calculus last semester.

OpenStudy (anonymous):

y=+or- sqrtx , y=x/2 to make it simpler

OpenStudy (anonymous):

come on guys, this isnt that hard.

OpenStudy (anonymous):

wait for a while, i'm solving it.

OpenStudy (anonymous):

what is really the given? it is lacking.

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

"find the volume of the solid obtained by rotating the regions bounded by the curves about the specified line.

OpenStudy (anonymous):

you have to find the point where the two curves meet, and then solve the integral from 0 to that point

OpenStudy (anonymous):

they meet at (0,0)

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

thats one yes, and then they cross again

OpenStudy (anonymous):

the answer is

OpenStudy (anonymous):

8pi over 3

OpenStudy (anonymous):

where do the two curves cross?

OpenStudy (anonymous):

the limits are 0 and 2.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

just substitute x=y^2 to x=2y that will become, y^2-2y=0 y(y-2)=0 y=0 y=2

OpenStudy (anonymous):

okay, let me double check now

OpenStudy (anonymous):

i had a mistake, the answer must be 10pi/3

OpenStudy (anonymous):

almost have it

OpenStudy (anonymous):

okay, i have the solution but i have a hard time encoding it.

OpenStudy (anonymous):

whats (2^1.5)/(2/3) -1

OpenStudy (anonymous):

i dont have my calculator :/

OpenStudy (anonymous):

3.242640687

OpenStudy (anonymous):

okay thats what i got... where did you get pi from?

OpenStudy (anonymous):

im looking in the back of the book for the answer to see whos right.

OpenStudy (anonymous):

2pi integral of (x^1/2-x/2)dx 2pi (x^3/2)/(3/2)-x^2/4, substitute x=2 2pi(2^3/2)(2/3)-2^2/4 2pi(4(2/3)-1) 2pi(8/3 - 1) 2 pi (5/3) (10pi)/3

OpenStudy (anonymous):

okay i did it waaaay wrong haha

OpenStudy (anonymous):

back of the book says 64pi/15

OpenStudy (anonymous):

is it correct now?

OpenStudy (anonymous):

is it the answer on the book?

OpenStudy (anonymous):

64pi/15, so its close

OpenStudy (anonymous):

okay im too tired for this, thanks for the help

OpenStudy (anonymous):

you are welcome, actually i can explain it, i think i just miss some terms in my solution. by the way, do you have skype account?